#30146 Home werk 20 series solution Polos Giveno y_y=0
Answer #1

Please see attached solution

Please like itQ=,hich hel Tero (6².-))) AE of me =(1-4) 1 2 o su l=u Y = 4 +çek Y= qtçet Please like it